# latex‘!h’ float specifier changed to ‘!ht’

### v1

@@ B.2 Associated weights for $T^{L}$
$T^{L}=\tilde{D}^{1/2}S^{L}\tilde{D}^{-1/2}.$ (14)
For any $i,j=1,\ldots,n$, let $t^{L}_{ij}$ represent the entry of the matrix $T^{L}$ in the $i$-th row and $j$-th column. Building on (14) and the formula (13) for the entries of $S^{L}$, we can derive that
@@ C.1 Proof of Proposition 1
$\underbrace{R,\cdots,R,}_{\text{($L-k$)-times}}\underbrace{F,\cdots,F,}_{\text{$r$-times}}\underbrace{B,\cdots,B,}_{\text{$(k-r)/2$-times}}\underbrace{F,\cdots,F}_{\text{$(k-r)/2$-times}}.$ (15)
@@
$\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{1}$-times}}\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{2}$-times}}\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{3}$-times}}\ldots,\underbrace{B,\cdots,B}_{\text{$q_{m}$-times}},\ldots,$ (16)
with $q_{1},\ldots,q_{m}\geq 1$ and $q_{1}+q_{2}+\ldots+q_{m}=(k-r)/2$. The number of paths of the form (16) in the augmented graph is hence bounded by $d^{q_{1}}d^{q_{2}}\ldots d^{q_{m}}=d^{(k-r)/2}.$ In a path we can in each step choose one of the three options $F,B,R.$ Thus there are at most $3^{L}$ different ways in (16) and for any $0\leq r\leq L$, $j\in{\mathcal{N}}^{r}(i)\setminus{\mathcal{N}}^{r-1}(i)$, and $k\in\mathcal{S}(r)$,
$\left|\mathcal{P}_{L}^{k,r}(i\to j)\right|\leq 3^{L}d^{(k-r)/2}.$ (17)
@@
$|{\mathcal{N}}^{r}(i)\setminus{\mathcal{N}}^{r-1}(i)|=d^{r},$ (18)
with the convention that ${\mathcal{N}}^{0}(i)=\{i\}$ and ${\mathcal{N}}^{-1}(i)=\emptyset$. Therefore, using (17) and (18),
@@
$\leq 3^{2L}(d+1)^{-2L}\sum_{r=0}^{L}\left(\sum_{k\in\mathcal{S}(r)}d^{\frac{k}{2}}\right)^{2}.$ (19)
@@ C.2 Proof of Proposition 2
$\omega_{i}(\ell_{1},\ldots,\ell_{L-1})=\frac{1}{(d_{i}+1)(d_{\ell_{1}}+1)\cdot\ldots\cdot(d_{\ell_{L-1}}+1)}\geq\frac{4^{1-L}}{d+1}.$ (20)
A direct calculation using (20) gives the result,
@@ Appendix D ADDITIONAL PROPOSITIONS AND PROOFS FOR $P=T^{L}$
$\widetilde{\omega}_{ij}(\ell_{1},\ldots,\ell_{L-1}):=\Big{(}\frac{d_{i}+1}{d_{j}+1}\Big{)}^{1/2}\omega_{i}(\ell_{1},\ldots,\ell_{L-1}).$ (21)
@@
Considering the single path $(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)$ with $d_{\ell_{1}},\ldots,d_{\ell_{L-1}}\leq 3$ and $d_{i}=d$, we obtain by applying (21) that
@@
Following the proof of Proposition 3, in which the nodes on the cycle $\mathcal{C}$ have been reindexed from 1 to $r$, starting from $i=\mathcal{C}\cap\mathcal{H}$, there exists at least one node $j\in\{2,\ldots,r\}$ such that $|\mathcal{P}_{L}(1\to j)|\geq 2^{L}/(r-1).$ Using (21), we can derive that
@@ E.1 Real data
In this subsection, we present some of the properties of the regression datasets used in the main text of this paper. Figure 5 shows the differences in terms of graph topologies and graph signal (assessed in terms of their “roughness”). In particular, we note that the county graphs are extremely regular in structure, while the twitch dataset exhibits a heavier tail in terms of degree distribution. Amongst the county graphs, the county election data exhibits the most smoothness. Interestingly, it is in this setting that the GCN convolution exhibits a significant advantage over the GraphSage convolution in both the denoising (see Figure 6) and prediction settings.
Figure 5: Properties of the datasets used in the experiment section.
Figure 6: Results for the denoising experiment on real data for the different regression datasets.
@@ E.2 Simulated Data
In this subsection, we provide examples of the different graphs generated to perform our synthetic experiments. The following set of 4 figures (Figures 7, 8, 9, and 10) highlight the different graph topologies, as well as the graph signal used.
@@
Figure 7: Barbell graph. Figure (a) shows the graph with corresponding graph signal $f^{*}=\cos(U\beta),\beta=(-\alpha,\alpha)$ with $\alpha=10$. Here $U$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.0001$ and figure (c) for $\alpha=10$.
@@
Figure 8: Binary Tree. Figure (a) shows the graph with corresponding graph signal $f^{*}=\cos(U\beta),\beta=(-\alpha,\alpha)$ with $\alpha=10$. Here $U$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.0001$ and figure (c) for $\alpha=10$.
@@
Figure 9: Power Law Cluster graph. Figure (a) shows the graph with corresponding graph signal $f^{*}=\cos(U\beta),\beta=(-\alpha,\alpha)$ with $\alpha=10$. Here $U$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.0001$ and figure (c) for $\alpha=10$.
@@
Figure 10: Latent Variable graph. Figure (a) shows the graph with corresponding graph signal $f^{*}=\cos(U\beta),\beta=(-\alpha,\alpha)$ with $\alpha=10$. Here $U$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.0001$ and figure (c) for $\alpha=10$.
@@
The next four sets of plots (Figures 11, 12, 13 and 14) highlight the bias-variance trade-off across different types of topologies. Overall, we observe a slower decay of the variance for the GraphSage convolution compared to the GCN convolution, for similar levels of bias.
@@
Figure 11: Bias-Variance as a function of $L$ for the Barbell graph, $m=20$.
@@
Figure 12: Bias-Variance as a function of $L$ for the tree (of degree 4).
@@
Figure 13: Bias-Variance as a function of $L$ for the power-law cluster graph ($m=1,p=0.05$).
@@
Figure 14: Bias-Variance as a function of $L$ for the sparsified latent variable graph ($p=0.75$).

### v2

@@ B.2 Associated Weights for $T^{L}$
$T^{L}=\tilde{D}^{1/2}S^{L}\tilde{D}^{-1/2}.$ (A.14)
For any $i,j=1,\ldots,n$, let $t^{L}_{ij}$ represent the entry of the matrix $T^{L}$ in the $i$-th row and $j$-th column. Building on (A.14) and the formula (A.13) for the entries of $S^{L}$, we can derive that
@@ C.1 Proof of Proposition 1
$\underbrace{R,\cdots,R,}_{\text{($L-k$)-times}}\underbrace{F,\cdots,F,}_{\text{$r$-times}}\underbrace{B,\cdots,B,}_{\text{$(k-r)/2$-times}}\underbrace{F,\cdots,F}_{\text{$(k-r)/2$-times}}.$ (A.15)
@@
$\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{1}$-times}}\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{2}$-times}}\ldots,\underbrace{B,\cdots,B,}_{\text{$q_{3}$-times}}\ldots,\underbrace{B,\cdots,B}_{\text{$q_{m}$-times}},\ldots,$ (A.16)
with $q_{1},\ldots,q_{m}\geq 1$ and $q_{1}+q_{2}+\ldots+q_{m}=(k-r)/2$. The number of paths of the form (A.16) in the augmented graph is hence bounded by $d^{q_{1}}d^{q_{2}}\ldots d^{q_{m}}=d^{(k-r)/2}.$ In a path we can in each step choose one of the three options $F,B,R.$ Thus there are at most $3^{L}$ different ways in (A.16) and for any $0\leq r\leq L$, $j\in{\mathcal{N}}^{r}(i)\setminus{\mathcal{N}}^{r-1}(i)$, and $k\in\mathcal{S}(r)$,
$\left|\mathcal{P}_{L}^{k,r}(i\to j)\right|\leq 3^{L}d^{(k-r)/2}.$ (A.17)
@@
$|{\mathcal{N}}^{r}(i)\setminus{\mathcal{N}}^{r-1}(i)|=d^{r},$ (A.18)
with the convention that ${\mathcal{N}}^{0}(i)=\{i\}$ and ${\mathcal{N}}^{-1}(i)=\emptyset$. Therefore, using (A.17) and (A.18),
@@
$\leq 3^{2L}(d+1)^{-2L}\sum_{r=0}^{L}\left(\sum_{k\in\mathcal{S}(r)}d^{\frac{k}{2}}\right)^{2}.$ (A.19)
@@ C.2 Proof of Proposition 2
$\omega_{i}(\ell_{1},\ldots,\ell_{L-1})=\frac{1}{(d_{i}+1)(d_{\ell_{1}}+1)\cdot\ldots\cdot(d_{\ell_{L-1}}+1)}\geq\frac{4^{1-L}}{d+1}.$ (A.20)
A direct calculation using (A.20) gives the result,
@@ Appendix D ADDITIONAL PROPOSITIONS AND PROOFS FOR $P=T^{L}$
$\widetilde{\omega}_{ij}(\ell_{1},\ldots,\ell_{L-1}):=\Big{(}\frac{d_{i}+1}{d_{j}+1}\Big{)}^{1/2}\omega_{i}(\ell_{1},\ldots,\ell_{L-1})$ (A.21)
@@
Considering the single path $(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)$ with $d_{\ell_{1}},\ldots,d_{\ell_{L-1}}\leq 3$ and $d_{i}=d$, we obtain by applying (A.21) that
@@
Following the proof of Proposition 3, in which the nodes on the cycle $\mathcal{C}$ have been reindexed from 1 to $r$, starting from $i=\mathcal{C}\cap\mathcal{H}$, there exists at least one node $j\in\{2,\ldots,r\}$ such that $|\mathcal{P}_{L}(1\to j)|\geq 2^{L}/(r-1).$ Using (A.21), we can derive that
@@ E.1 Real Data
In this subsection, we present some of the properties of the datasets used in the main text of this paper. Figure 5 shows the differences in terms of graph topologies and graph signal (assessed in terms of their “roughness”). In particular, we note that the county graphs are extremely regular in structure, while the Twitch dataset exhibits a heavier tail in terms of degree distribution. Among the county graphs, the county election data is the smoothest. Interestingly, in this setting, the GCN convolution outperforms the GraphSAGE convolution in both denoising (see Figure 6) and prediction.
Figure 5: Properties of the datasets used in the real-data experiment section. The last 3 columns measure the signal roughness, defined either in terms of the $\ell_{2}$ norm ($\|\Delta{\bf Y}\|^{2}_{2}/|\mathcal{E}|=[\sum_{(i,j)\in\mathcal{E}}(Y_{i}-Y_{j})^{2}]/|\mathcal{E}|$) or the $\ell_{\infty}$ norm ($\|\Delta{\bf Y}\|_{\max}=\max_{(i,j)\in\mathcal{E}}|Y_{i}-Y_{j}|$). The last column represent the graph smoothness over the neighborhood (after one convolution $S$).
Figure 6: Results of the denoising experiment on real data across different regression datasets.
@@ E.2 Simulated Data
In this subsection, we provide examples of the different graphs used in our synthetic experiments. The following four figures (Figures 7, 8, 9, and 10) illustrate the various graph topologies and the corresponding graph signals.
@@
Figure 7: Barbell graph. Figure (a) shows the graph with corresponding graph signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta),\beta=(-\alpha,\alpha)^{\top}$ with $\alpha=10$. Here $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.001$ and figure (c) for $\alpha=10$.
@@
Figure 8: Binary Tree. Figure (a) shows the graph with corresponding graph signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta),\beta=(-\alpha,\alpha)^{\top}$ with $\alpha=10$. Here $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.001$ and figure (c) for $\alpha=10$.
@@
Figure 9: Power Law Cluster graph. Figure (a) shows the graph with corresponding graph signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta),\beta=(-\alpha,\alpha)^{\top}$ with $\alpha=10$. Here $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.001$ and figure (c) for $\alpha=10$.
@@
Figure 10: Latent Variable graph. Figure (a) shows the graph with corresponding graph signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta),\beta=(-\alpha,\alpha)^{\top}$ with $\alpha=10$. Here $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ represents the two dimensional spectral embedding of the graph. Figure (b) shows the spectral embedding, colored by signal for $\alpha=0.001$ and figure (c) for $\alpha=10$.
The next set of plots (Figures 11, 12 and 13) highlights the optimal number of convolutions as a function of the graph roughness. Overall, we observe the same phenomenon on the GCN and GraphSAGE convolutions as for the latent variable graph described in the main text: as the roughness of the graph increases, the optimal number of convolutions decreases. In all cases, we observe that when the neighborhood becomes too uninformative, the optimal number of layers increases again. This, however, needs to be understood alongside with the error, which increases substantially.
Figure 11: Optimal number of convolutions as a function of the roughness of $\mathbf{f}^{*}$ (here defined as $\|\Delta{\bf f^{*}}\|_{2}=\sqrt{[\sum_{(i,j)\in\mathcal{E}}(f_{i}^{*}-f_{j}^{*})^{2}]/|\mathcal{E}|}$) on the Barbell graph. Each column corresponds to a different clique size $m$, (with, for instance, 10 meaning that the two cliques on either side of the Barbell graph are of size 10) and each row to a value of the noise $\sigma^{2}$.
Figure 12: Optimal number of convolutions as a function of the roughness of $\mathbf{f}^{*}$ (here defined as $\|\Delta{\bf f^{*}}\|_{2}=\sqrt{[\sum_{(i,j)\in\mathcal{E}}(f_{i}^{*}-f_{j}^{*})^{2}]/|\mathcal{E}|}$) on a tree graph. Each column corresponds to a different branching number (or number of leaves) $k$, and each row to a value of the noise $\sigma^{2}$: the top row corresponds to a value of $\sigma^{2}=0.5$, while the bottom row corresponds to $\sigma^{2}=2$.
Figure 13: Optimal number of convolutions as a function of the roughness of $\mathbf{f}^{*}$ (here defined as $\|\Delta{\bf f^{*}}\|_{2}=\sqrt{[\sum_{(i,j)\in\mathcal{E}}(f_{i}^{*}-f_{j}^{*})^{2}]/|\mathcal{E}|}$) on the power-law cluster graph. Each column corresponds to a different clustering parameter $p$, and each row to a value of the noise $\sigma^{2}$: the top row corresponds to a value of $\sigma^{2}=0.5$, while the bottom row corresponds to $\sigma^{2}=2$.
@@
The next sets of plots (Figures 14, 15, 16, 17, 18 and 19) highlight the bias-variance trade-off across different types of topologies. Overall, we observe a slower decay of the variance for the GraphSAGE convolution compared to the GCN convolution, for similar levels of bias.
@@
Figure 14: Bias-Variance as a function of $L$ for the Barbell graph, clique size $m=20$, $\sigma^{2}=2$.
@@
Figure 15: Bias-Variance as a function of $L$ for the tree (with branching number equal to 2), $\sigma^{2}=1$.
Figure 16: Bias-Variance as a function of $L$ for the tree (with branching number equal to 4), $\sigma^{2}=1$.
@@
Figure 17: Bias-Variance as a function of $L$ for the power law cluster graph ($m=1,p=0.1$), $\sigma^{2}=1$.
@@
Figure 18: Bias-Variance as a function of $L$ for the latent variable graph, $\sigma^{2}=1$.
Figure 19: Bias-Variance as a function of $L$ for the sparsified latent variable graph ($p=0.5)$, $\sigma^{2}=1$.
Figure 20: Variance decay of different estimators on a tree with a branching factor of 4 over four levels, with a clique of 10 nodes added (left) or a star of 10 nodes added (right).